OpenStudy (anonymous):

Medal reward!! A small publishing company is planning to publish a new book. The production costs will include one-time fixed costs (such as editing) and variable costs (such as printing). The one-time fixed costs will total $26,109. The variable costs will be $12 per book. The publisher will sell the finished product to bookstores at a price of $18.75 per book. How many books must the publisher produce and sell so that the production costs will equal the money from sales?

OpenStudy (nurali):

Let x = no. of books Cost = 12x + 26109 Revenue = 18.75x Rev = cost 18.75x = 12x + 26109 18.75x - 12x = 26109 6.75x = 26109 x = 26109/6.75 x = 3868 books need to be sold to cover production costs

OpenStudy (anonymous):

Thank you! :)

OpenStudy (nurali):

Anytime.